OpenStudy (anonymous):

Find the modulo class to which the number belongs for the given modulo system. 57, mod 12 A. 6 B. 9 C. 1 D. 3

OpenStudy (amistre64):

what is the remainder of 57/12?

OpenStudy (anonymous):

4 but thats what I dont really understand, 4 is not one of the choices

jimthompson5910 (jim_thompson5910):

Similar Example 55/12 = 4 remainder 7 The remainder is 7, which means that 55 (mod 12) = 7 (mod 12) This basically says that they have the same remainder when you divide them by 12

jimthompson5910 (jim_thompson5910):

So do the same for 57

OpenStudy (anonymous):

oh ok i get it, so the answer is 9

jimthompson5910 (jim_thompson5910):

yep 57/12 = 4 remainder 9 so 57 (mod 12) = 9 (mod 12)

OpenStudy (anonymous):

ok, thank u so much:)

OpenStudy (amistre64):

ida had that but the site crashed on me :/

OpenStudy (amistre64):

57/12 = (48 + 9)/12 = 4 + 9/12 so the remainder is not "4"